A continuous random variable X has the following pdf f (x) = kxe− 1 x ; k is a constant.x > 0

 X2

(a) Find the value of k for fX(x) to be a valid probability density function.

(b) Find the cumulant generating function of X.

(b) Using the cumulant generating function or otherwise, find the mean and variance of X.


1
Expert's answer
2021-11-22T19:54:48-0500

(a)


"\displaystyle\int_{-\infin}^{\infin}f(x)dx=\displaystyle\int_{0}^{\infin}kxe^{-x}dx"

"=k\lim\limits_{A\to \infin}\displaystyle\int_{0}^{A}xe^{-x}dx"

"=k\lim\limits_{A\to \infin}[-xe^{-x}]\begin{matrix}
 A \\
 0
\end{matrix}+k\lim\limits_{A\to \infin}\displaystyle\int_{0}^{A}e^{-x}dx"

"=0-k\lim\limits_{A\to \infin}[e^{-x}]\begin{matrix}
 A \\
 0
\end{matrix}"

"=k=1=>k=1""f(x) = \begin{cases}
 xe^{-x} &x>0 \\
 0 & x\leq0
\end{cases}"

(b)


"M_x(t)=E(e^{xt})=\displaystyle\int_{-\infin}^{\infin}e^{xt}f(x)dx"

"=\lim\limits_{A\to \infin}\displaystyle\int_{0}^{A}xe^{x(t-1)}dx"

"\int xe^{(t-1)x}dx=\dfrac{1}{t-1}xe^{(t-1)x}-\dfrac{1}{t-1}\int e^{(t-1)x}dx"




"=\dfrac{1}{t-1}xe^{(t-1)x}-\dfrac{1}{(t-1)^2}e^{(t-1)x}+C"


"M_x(t)" exists for "t<1"


"M_x(t)=\lim\limits_{A\to \infin}\displaystyle\int_{0}^{A}xe^{x(t-1)}dx"

"=\lim\limits_{A\to \infin}[\dfrac{1}{t-1}xe^{(t-1)x}-\dfrac{1}{(t-1)^2}e^{(t-1)x}]\begin{matrix}
 A \\
 0
\end{matrix}"

"=-0-0-(-0-\dfrac{1}{(t-1)^2})=\dfrac{1}{(t-1)^2}, t<1"

"K_x(t)=\ln(M_x(t))=-2\ln(1-t), t<1"

c)


"K_x'(t)=\dfrac{2}{1-t}"

"K_x''(t)=\dfrac{2}{(1-t)^2}"

"mean=E(X)=K_x'(0)=\dfrac{2}{1-0}=2"

"Var(X)=K_x''(0)=\dfrac{2}{(1-0)^2}=2"